marks People Window Help Manifest l Xfinity S x: M Join Alpha Epsilon : × × Ψ webwork xl weewerk SPtH work.uits.iu.edu/webwork2/SP19-BL-MATH-D116-2936/Section 1.3 A/4/key-ynDW6CpaioxSb2UmzNpObmbi MATHEMATICAL ASSOCIATION OF AMERICA webwork / sp19-bl-math-d116-2936/section 1.3 a/4 Section 1.3 A: Problem 4 Previous ProblemProblem List Next Problem (1 point) Let U be a universal set with subsets A and B such that nU)121, A) 3,(B)64, and mAnB) 22. Find n(AUB) n(A u B) = Submit Answers Preview My Answers You have attempted this problem 4 times. Your overall recorded score is 0% You have unimited attempts remaining. Email instructor 名0 F3 9 6 5 4 3 2
Answer #1

n(L)-121 n(A) = 53 (B) 64 n (An B) 22 n(A) n(A)-121 n(A) 121-53 68 n(B) + n(В)-121 n(B) = 121-64-57 n(AU B) = n(A) + n(B)-